OpenStudy (anonymous):

the first term of a sequence is 12 and each term is 7 less than the last. what is the 25th term?

OpenStudy (dumbcow):

arithmetic sequence \[a _{n} = a _{1} + d(n-1)\] a1 = 12 d = -7 n = 25

OpenStudy (anonymous):

thats what i was thought. thanks guys.

OpenStudy (anonymous):

Oh ... My bad. The correct answer is: n = 0, An = 12 n = 1, An = 12-7 n = 2, An = 12-7*2 ... n = 24, An = 12-7*24

OpenStudy (dumbcow):

diogo, usually with an arithmetic sequence the 1st term means n=1 Not n=0

OpenStudy (anonymous):

if you start at n = 0, then n = 25 is really the 26th term

OpenStudy (anonymous):

ah, nvm, you caught it. nice job :)